# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ 4.1. Grading
$\partial Q=0=[u^{\prime\prime}]-[u^{\prime}]$
in $H_{2}(Y^{0},L_{j_{1}}\cup L_{j_{2}}\cup\psi(\Phi_{H}(K_{j_{3}})))$ where $\psi$ is the Lagrangian isotopy induced by $\delta$ and so
$u^{\prime}]=[u^{\prime\prime}]$
where $u^{\prime\prime}$ is in a $T^{4}$-fiber bounded by the three linear Lagrangians. ∎
The proof for 2 cotangent fibers and 1 U-shape as in Figure 9 is similar, where $\delta(s,t)$ is a homotopy moving the right side of the triangle, $L_{j_{1}}$ across to $\epsilon^{\prime}$ and $W_{0}(\beta)$ covers again the shaded region. Now $\Phi_{H}$ is chosen to have $L_{j_{1}}$ match the perturbed $K_{j_{2}}$ and $K_{j_{3}}$. Then in a fiber the triangle is bounded by $\Phi_{\delta_{K_{j}}(0)\to\epsilon^{\prime}}(\ell_{j_{1}}\cup\ell_{j_{2}}\cup\ell_{j_{3}})$ from going from S-S to $\epsilon^{\prime}$ along the path for $W_{0}(K_{j})$ in Definition 3.6. We can apply the symplectomorphism $\Phi_{\delta_{K_{j}}(0)\to\epsilon^{\prime}}^{-1}$ to obtain a bijection between those triangles and the linear ones, that preserves their weights.
Lemma 4.3.
The compositions with two U-shaped Lagrangians and one cotangent fiber Lagrangian
$HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$
and with two cotangent fiber Lagrangians and one U-shaped Lagrangian
$HF(K_{j_{2}}[-1],K_{j_{3}}[-1])\otimes HF(L_{j_{1}},K_{j_{2}}[-1])\to HF(L_{j_{1}},K_{j_{3}}[-1]),$
after rescaling, are products in the abelian variety fiber.
Proof.
One cotangent fibered and two U-shaped: each Floer cochain complex is a direct sum of fiber cochain complexes over two base as can be seen in Figure 7. For example
(4.11) $CF(L_{j_{2}},K_{j_{3}}[-1])=CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\oplus CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$
in fibers over $b_{4}$ and $b_{5}$. The differential for each cochain complex goes from the intersection point on the right (degree 1-1) to that on the left (degree 0). For example, we have an exact sequence computing Floer cohomology from
(4.12) $0\to CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})[-1]\xrightarrow[]{\partial}CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\to HF(L_{j_{2}},K_{j_{3}}[-1])\to 0$
where $\partial$ counts images of bigon sections of $W_{0}$ over the bigon in the base between $b_{4}$ and $b_{5}$.
Lemma 4.2 showed that composition $M^{2}$ of the degree 0 components
(4.13) $M^{2}:CF^{0}(L_{j_{2}},K_{j_{3}}[-1])\otimes CF^{0}(L_{j_{1}},L_{j_{2}})$ $\to CF^{0}(L_{j_{1}},K_{j_{3}}[-1])$
$M^{2}:CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\otimes CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$ $\to CF((L_{j_{1}})_{b_{2}},(K_{j_{3}}[-1])_{b_{2}})$
can be computed in the fiber with linear Lagrangians. For degree reasons, given that $M^{2}$ is of degree 0, there are only two other possibilities for the triangles:
(4.14) $M^{2}:CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})\otimes CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})\to CF((L_{j_{1}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$
and
(4.15) $M^{2}:CF((L_{j_{2}})_{b_{4}},(K_{j_{3}}[-1])_{b_{4}})\otimes CF((L_{j_{1}})_{b_{1}},(L_{j_{2}})_{b_{1}})\to CF((L_{j_{1}})_{b_{3}},(K_{j_{3}}[-1])_{b_{3}}).$
These show that $M^{2}$ passes to cohomology, as follows. Recall that $M^{2}$ satisfies a Leibniz rule, in that modulo 2,
Applying this to Equation (4.14), we find that for $c\in CF((L_{j_{2}})_{b_{5}},(K_{j_{3}}[-1])_{b_{5}})$ and $c^{\prime}\in CF((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$
(4.17) $\partial\circ M^{2}(c,c^{\prime})=M^{2}(\partial(c),c^{\prime})+\cancel{M^{2}(c,\partial(c^{\prime}))}$
where the second term is zero because $\partial(c^{\prime})=0$ since there $\partial$ increase degree by 1, $c^{\prime}$ is degree 0, and there are no intersection points of degree 1. By the analogous calculation for Equation (4.15), we find that $M^{2}$ passes to the quotient by the image of $\partial$, to cohomology:
(4.18) $HF(L_{j_{2}},K_{j_{3}}[-1])\otimes HF(L_{j_{1}},L_{j_{2}})\to HF(L_{j_{1}},K_{j_{3}}[-1])$
and it can be computed with representatives in the degree 0 components, whose product can be reduced to products in the abelian variety fiber. A similar argument can be done in the second case.
To show that we can make the products equal that of the fiber products by rescaling the intersection points, first consider the first case. There, the Lagrangian $K_{j_{3}}$ is moved, which moves $b_{2}$ (output of $M^{2}$) and $b_{4}$ (an input of $M^{2}$). Pick intersection points over these base points, $p_{1}\in Y^{0}_{b_{4}}$ the fiber over $b_{4}$, and $q\in Y^{0}_{b_{2}}$. Then
$\rho(u)\rho(u^{\prime})^{-1}=\rho(u)\rho(u^{\prime\prime})^{-1}=e^{2\pi if(p_{1})}e^{-2\pi if(q)}$
by the last equation in the proof of [ACLL23b, Corollary 2.6] for a certain function $f$ defined there and depending on the Lagrangian isotopy. In other words, the original triangle and the fiber triangle differ an amount that only depends on the intersection points $p_{1},q$. Thus rescaling $p_{1}\mapsto e^{2\pi if(p_{1})}p_{1}$ and $q\mapsto e^{2\pi if(q)}q$, we find that for $p_{0}\in CF^{0}((L_{j_{1}})_{b_{0}},(L_{j_{2}})_{b_{0}})$
(4.19) $p_{1}\cdot p_{0}=\sum_{q\in L_{j_{1}}\cap K_{j_{3}}}C_{p_{0}p_{1}q}q$
where $C_{p_{0}p_{1}q}$ is the count of triangles between linear Lagrangians in a fiber. The proof for the second case is similar. ∎
5. Proof of main theorem
Now we prove the main theorem.
Theorem 5.1.
The full subcategories $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ of the Fukaya category $H^{*}FS(Y^{0},W_{0})$ are equivalent to $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ and $D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, respectively. Furthermore, for the full subcategory $\mathcal{F}:=\mathcal{F}_{1}\cup\mathcal{F}_{2}$, there is an equivalence
(5.1) $\mathcal{F}\xrightarrow[]{\cong}D^{b}_{\mathcal{L}}\mathrm{Coh}(X).$
Theorem 5.2 ([Can20, §6]).
There exists a commutative diagram as follows.
$D^{b}_{\mathcal{L}}\mathrm{Coh}(V)$$D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$$H^{0}\mathrm{Fuk}(V^{\vee})$$H^{0}\mathrm{FS}(Y,-v_{0})$$\scriptstyle{\iota^{*}}$$\scriptstyle{\text{HMS on }V}$$\scriptstyle{\text{HMS on }H=\Sigma_{2}}\scriptstyle{\cup}$
Corollary 5.3 (HMS on $\mathcal{F}_{1}$ factor).
$\mathcal{F}_{1}\cong D^{b}_{\mathcal{L}}\mathrm{Coh}(H)$ under $L_{k}\mapsto\mathcal{L}^{k}|_{H}$.
Proof.
By the right vertical arrow in Theorem 5.2, $D^{b}_{\mathcal{L}}\mathrm{Coh}(H)\cong\mathcal{F}_{1}^{\prime}$ for $\mathcal{F}_{1}^{\prime}\subset H^{0}\text{FS}(Y,-v_{0})$ the set of U-shape Lagrangians in $v_{0}:Y\to\mathbb{C}$. Then $\mathcal{F}_{1}^{\prime}\cong\mathcal{F}_{1}$ under the affine transformation on the base from $v_{0}$ to $W_{0}=T^{\epsilon}(v_{0}-1)$. ∎
Corollary 5.4 (HMS on $\mathcal{F}_{2}$ factor).
$\mathcal{F}_{2}\cong D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$ under $K_{j}\mapsto\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}$.
Proof.
This follows from Lemma 3.25, HMS for abelian varieties [Fuk02, ACLL], the Künneth formula for Ext, and that $\text{Ext}(\mathcal{O}_{\mathbb{C}},\mathcal{O}_{\mathbb{C}})=\mathbb{C}[x]$. ∎
@@ 5. Proof of main theorem
$\mathcal{F}=\langle\mathcal{F}_{1},\mathcal{F}_{2}\rangle\to D^{b}_{\mathcal{L}}\mathrm{Coh}(X)=\langle j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1)),\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})\rangle$
defined on $\mathcal{F}_{1}$ by
(5.2) $L_{k}\mapsto j_{*}(p^{*}\mathcal{L}^{\otimes k}|_{H}\otimes\mathcal{O}_{p}(-1))$
and on $\mathcal{F}_{2}$ by
(5.3) $K_{j}\mapsto\pi^{*}(\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}).$
is an equivalence within each factor of the semi-orthogonal decomposition by Corollaries 5.3 and 5.4. Since $j:E\to X$ is a closed immersion,
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{1}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),j_{*}(p^{*}i^{*}(\mathcal{L}^{k_{2}}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)))=\operatorname{Ext}(\mathcal{L}^{k_{1}}|_{H},\mathcal{L}^{k_{2}}|_{H}).$
For morphisms between factors it is zero on both sides in one direction by Equation (2.1) on the B-side and Lemma 3.24 on the A-side. That is,
$\operatorname{Ext}(\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}}),j_{*}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{p}(-1)))=0=HF(K_{j},L_{k}).$
In the other direction, the equivalence on morphisms reduces to HMS for $H$, as follows. By Equation (2.3)
$\operatorname{Ext}(j_{*}(p^{*}i^{*}(\mathcal{L}^{k}\boxtimes\mathcal{O}_{\mathbb{C}})\otimes\mathcal{O}_{E}(E)),\pi^{*}(\mathcal{L}^{j}\boxtimes\mathcal{O}_{\mathbb{C}}))\cong\operatorname{Ext}(\mathcal{L}^{k}|_{H},\mathcal{L}^{j}|_{H})[-1].$
@@
$HF(L_{k},L_{j})[-1]\cong HF(L_{k},K_{j})$
because the grading on $W_{0}(K_{i})$ is $\alpha_{i}=0$ which points in the opposite direction from $W_{0}(L_{i})$ constructed by extending as in Figure 4.
To show it’s a functor it remains to prove it respects composition between the two components. This follows from Lemma 4.3 and HMS for abelian varieties [ACLL, Fuk02].

### v2

@@ 4. Proof of main theorem
(4.21) $\operatorname{Ext}^{*}(\text{pr}_{V}^{*}\mathcal{L}^{\otimes i}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}},\text{pr}_{V}^{*}\mathcal{L}^{\otimes j}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}})\cong HF^{*}(K_{i},K_{j}).$
@@
$\mathcal{A}_{L,K}=\langle\mathcal{A}_{L},\mathcal{A}_{K}\rangle\to D^{b}_{\mathcal{L}}\mathrm{Coh}(X)=\langle j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1)),\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})\rangle$
defined by
(4.22) $\mathcal{A}_{L}\ni L_{j_{1}}$ $\mapsto j_{*}(p^{*}\mathcal{L}^{\otimes{j_{1}}}|_{H}\otimes\mathcal{O}_{p}(-1))\in j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1))$
$\mathcal{A}_{K}\ni K_{j_{2}}$ $\mapsto\pi^{*}(\text{pr}_{V}^{*}\mathcal{L}^{\otimes j_{2}}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}})\in\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$
is an equivalence within each factor of the semi-orthogonal decomposition by Corollaries 4.3 and 4.4. That is, letting
(4.23) $\mathcal{L}^{j_{1},H}:=j_{*}(p^{*}\mathcal{L}^{\otimes{j_{1}}}|_{H}\otimes\mathcal{O}_{p}(-1)),\qquad\mathcal{L}^{j_{2},V\times\mathbb{C}}:=\pi^{*}(\text{pr}_{V}^{*}\mathcal{L}^{\otimes j_{2}}\otimes\text{pr}_{\mathbb{C}}^{*}\mathcal{O}_{\mathbb{C}}),$
then since $j:E\to X$ is a closed immersion and $\otimes\mathcal{O}_{p}(-1)$ is an auto-equivalence,
(4.24) $\operatorname{Ext}^{*}(\mathcal{L}^{j_{1},H},\mathcal{L}^{j_{2},H})=\operatorname{Ext}^{*}(\mathcal{L}^{j_{1}}|_{H},\mathcal{L}^{j_{2}}|_{H})=HF^{*}(L_{j_{1}},L_{j_{2}}),$
(4.25) $\operatorname{Ext}^{*}(\mathcal{L}^{j_{1},V\times\mathbb{C}},\mathcal{L}^{j_{2},V\times\mathbb{C}})=HF^{*}(K_{j_{1}},K_{j_{2}}),$
and composition is respected within each factor.
For morphisms between factors, it is zero on both sides in one direction, by Equation (2.1) on the B-side and Lemma 3.28 on the A-side. That is,
$\operatorname{Ext}(\mathcal{L}^{j_{1},V\times\mathbb{C}},\mathcal{L}^{j_{2},H})=0=HF(K_{j_{1}},L_{j_{2}}).$
In the other direction, the equivalence on morphisms reduces to HMS for $H$, as follows. By Equations (2.2) and (2.3)
(4.26) $\operatorname{Ext}(\mathcal{L}^{j_{2},H},\mathcal{L}^{j_{1},V\times\mathbb{C}})\cong\operatorname{Ext}(\mathcal{L}^{j_{2}}|_{H},\mathcal{L}^{j_{1}}|_{H})[-1].$
@@
(4.27) $HF(L_{j_{2}},L_{j_{1}})[-1]\cong HF(L_{j_{2}},K_{j_{1}}).$
Lastly, composition is respected between the two components by Lemma 3.31 and HMS for abelian varieties [ACLL_tori, Proposition 4.18], [fuk_abel]. That is, the following two diagrams commute:
$HF^{d_{2}}(L_{j_{2}},K_{j_{3}}[1])\otimes HF^{d_{1}}(L_{j_{1}},L_{j_{2}})$$HF^{d_{1}+d_{2}}(L_{j_{1}},K_{j_{3}}[1])$$\operatorname{Ext}^{d_{2}}(\mathcal{L}^{j_{2}}|_{H},\mathcal{L}^{j_{3}}|_{H})\otimes\operatorname{Ext}^{d_{1}}(\mathcal{L}^{j_{1}}|_{H},\mathcal{L}^{j_{2}}|_{H})$$\operatorname{Ext}^{d_{1}+d_{2}}(\mathcal{L}^{j_{1}}|_{H},\mathcal{L}^{j_{3}}|_{H})$$\scriptstyle{M^{2}}\scriptstyle{\cong}\scriptstyle{\cong}\scriptstyle{\otimes}$
and
[…]